# Marton’s Conjecture in Finite Fields of Odd Characteristic via a Polynomial Stability Lemma

### v1

@@ 4 Lemma S1: Polynomial Chang span bound
Under the doubling assumption $|A+A|\leq K|A|$, we have $\alpha\geq 1/K$, so $\log(1/\alpha)\leq\log K$. Choosing $\tau=K^{-c_{0}}$ for a suitably small $c_{0}>0$, we obtain
@@ 6 Lemma S2: Dispersion and polynomial improvement
Plancherel gives $\mathsf{E}(A^{\prime})=|G^{\prime}|\sum_{\psi}|\widehat{\mathbf{1}_{A^{\prime}}}(\psi)|^{4}$. The trivial character contributes $\alpha^{\prime 4}|G^{\prime}|$. If there is a nontrivial $\chi$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq\eta\alpha^{\prime}$, the sum gains at least $\eta^{4}\alpha^{\prime 4}$ from that term alone, hence (4) with $c_{*}=1$. The inequality $|A^{\prime}+A^{\prime}|\leq|A^{\prime}|^{4}/\mathsf{E}(A^{\prime})$ then yields the claimed decrement. ∎
@@ A.1 Quantitative BSG and covering
Apply the $L^{4}$-energy form of BSG (see Tao–Vu [3, Thm. 2.29]; cf. [6, 5]). Replacing the usual parameter $K^{\prime}$ by $K^{\prime}=(1+\beta)^{-1}$ yields the stated polynomial bounds in $\beta$. ∎
@@
Ruzsa’s covering plus Plünnecke inequalities; see [3, Ch. 2]. The last assertion is a reparametrization giving an explicit decrement. ∎
@@ Appendix E Standing assumptions and parameter window
Throughout we fix absolute constants $c_{0},c\in(0,1)$ and apply Rudin–Chang in the regime $\alpha\in[K^{-1},1/2]$. All constants ($C_{1},C_{2},C$) are independent of $K$ and $p$; exponents can be traced explicitly by following the proofs of Lemmas 4.1,6.4,5.1.

### v2

@@ 1.1 Context and related work
Concurrent and subsequent work.
We note that, subsequent to the initial development of this work, Gowers, Green, Manners, and Tao [2] established Marton’s conjecture for abelian groups of bounded torsion using entropy-theoretic methods. Since finite fields are groups of bounded torsion, their result also covers finite fields of odd characteristic. The present work provides an independent, direct Fourier-analytic proof with explicit spectral stability dichotomies and quantitative polynomial bounds in the finite-field setting.
@@ 4 Lemma S1: Polynomial Chang span bound
By the standing assumptions (Appendix E) and the modeling Lemma A.1, we may, without loss of generality, work in a Freiman model where the density satisfies $\alpha\in[K^{-C},1/2]$. In particular $\log(1/\alpha)\ll\log K$. Choosing $\tau=K^{-c_{0}}$ for a suitably small $c_{0}>0$, we obtain
@@ 6 Lemma S2: Dispersion and polynomial improvement
Plancherel gives $\mathsf{E}(A^{\prime})=|G^{\prime}|\sum_{\psi}|\widehat{\mathbf{1}_{A^{\prime}}}(\psi)|^{4}$. The trivial character contributes $\alpha^{\prime 4}|G^{\prime}|$. If there is a nontrivial $\chi$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq\eta\alpha^{\prime}$, the sum gains at least $\eta^{4}\alpha^{\prime 4}$ from that term alone, hence (4) with $c_{*}=1$. The inequality $|A^{\prime}|^{4}\leq\mathsf{E}(A^{\prime})\,|A^{\prime}+A^{\prime}|$ then yields the claimed decrement. ∎
@@ Appendix A Appendix A: Toolbox — BSG, covering, and constant ledger
Lemma A.1 (Density window via Freiman modeling).
Let $A\subset G$ with $|A+A|\leq K|A|$. Then there exists a finite abelian group $G_{0}$ with
$|G_{0}|\ \leq\ K^{C}\,|A|$
and a Freiman isomorphism $\phi:A\to A_{0}\subset G_{0}$ such that $\alpha_{0}:=|A_{0}|/|G_{0}|\geq K^{-C}$. Moreover, passing to $G_{0}$ preserves all additive-combinatorial inequalities used in this paper up to absolute constants, so any estimate proved under the standing assumption $\alpha\in[K^{-C},1/2]$ applies to $A_{0}$ and then pulls back to $A$.
Proof.
This is a standard consequence of Ruzsa’s modeling lemma; see [4, Thm. 2.29] and [5]. The bound $|G_{0}|\leq K^{C}|A|$ is polynomial in $K$, hence $\alpha_{0}\geq|A|/|G_{0}|\geq K^{-C}$. Freiman isomorphism preserves sumset relations and energies appearing in our arguments. ∎
@@ A.1 Quantitative BSG and covering
Apply the $L^{4}$-energy form of BSG (see Tao–Vu [4, Thm. 2.29]; cf. [7, 6]). Replacing the usual parameter $K^{\prime}$ by $K^{\prime}=(1+\beta)^{-1}$ yields the stated polynomial bounds in $\beta$. ∎
@@
Ruzsa’s covering plus Plünnecke inequalities; see [4, Ch. 2]. The last assertion is a reparametrization giving an explicit decrement. ∎
@@ Appendix E Standing assumptions and parameter window
Throughout we fix absolute constants $c_{0},c\in(0,1)$ and apply Rudin–Chang in the regime $\alpha\in[K^{-C},1/2]$ (achieved by Lemma A.1). All constants ($C_{1},C_{2},C$) are independent of $K$ and $p$; exponents can be traced explicitly by following the proofs of Lemmas 4.1,6.4,5.1.